OpenStudy (anonymous):

Write the expression in standard form. 6/4-13i

OpenStudy (midhun.madhu1987):

Multiply the numerator and denominator by the conjugate 4+13i and simplify

OpenStudy (anonymous):

i did and it didn add up to any of he answers

OpenStudy (anonymous):

@midhun.madhu1987 can you show me your work ?

geerky42 (geerky42):

\[\dfrac{6}{4-13i}\Longrightarrow\dfrac{6}{4-13i}\cdot\dfrac{4+13i}{4+13i}\Longrightarrow\dfrac{6(4-13i)}{(4)^2-(13i)^2}\]I'm sure you can handle the rest.

OpenStudy (midhun.madhu1987):